if the speed of light in air 3 x 10 ⁸ m/s find the speed of light in diamond of refraction index 2.419

1
Expert's answer
2022-01-30T13:51:43-0500

Answer


speed of light in air c=3 x 10 ⁸ m/s

refraction index n=2.419

Speed of light in diamond

"v=\frac{c}{n}\\=\frac{3 *10 ⁸}{2.419}\\=1.24*10 ⁸m/s"